# Prediction of linear fractional stable motions using codifference

### v1

@@ 1 Introduction
One can consider several alternatives to the covariance to build forecasting methods. First, though moments of order $p$ are infinite when $p\leq\alpha$, conditional moments can be finite for higher orders [60, Chapter 5], but it is under a set of assumptions which does not hold when future increments are obtained by adding independent variables to past observations, like in the LFSM [60, Theorem 5.1.3]. The conditional expectation is however finite when another restrictive condition is met, namely when $\alpha>1$ [60, Theorem 5.2.2]. The obtained analytic formula can however not be generalized for a conditioning on a number of lagged increments larger than 1 [59, 35]. A second alternative would consist in describing the dependence structure with copulas. It is however not convenient because it requires an expression for the cumulative distribution function, which, in the case of stable variables, can only be obtained numerically. A third alternative would be to depict the dependence by the mean of the multivariate characteristic function instead of the multivariate cumulative distribution function. Indeed, the spectral measure, which is derived from this multivariate characteristic function, entirely describes the dependence between stable variables. Since it is a function, the spectral measure is an object of infinite dimension, which is thus difficult to estimate from limited observations. Instead, one often considers the codifference [40, 7, 56, 72], which simplifies the dependence structure contained in the spectral measure as the correlation or Kendall rank correlation do with the copula.
In this article, similarly to the Cholesky decomposition of a Gaussian vector, we propose a decomposition of discrete-time observations of an LFSM in independent stable variables. The decomposition has not exactly the same dependence structure as an LFSM but it has the same codifference. From this decomposition, we can propose a forecasting method. The evaluation of this method, first theoretically then on simulations, shows a new way of interpreting the parameters of an LFSM, namely the stability parameter $\alpha$ and the Hurst exponent $H$. Indeed, in addition to the traditional persistent, independent, and antipersistent cases, a new property of the time series appears for very small values of $\alpha$. The method is also compared to a previous study on the fBm using real time series of volatilities or of FX rates. The performance of our forecasting method in this real framework is promising and legitimises the model.
@@
The rest of the article is organized as follows. Section 2 presents the model along with some of its properties and the concept of codifference. In Section 3, we introduce the decomposition of a vector of observations of the LFSM and deduce the forecasting method. A simulation study and an application to real data are proposed respectively in Sections 4 and 5. Section 6 concludes.
2 The LFSM
@@
In this section, we first introduce LFSM along with codifference, then we briefly present simulation and estimation methods.
2.1 Definitions and properties
Stable distributions have been explored in pioneering works of Lévy and Khinchine, about a century ago. Several equivalent definitions of this kind of distribution exist [60]. Unfortunately, no analytic expression is available for the probability density function (pdf) but we know the characteristic function, which is enough to define this distribution and from which we can get the pdf numerically thanks to an inverse Fourier transform [57, 5].
Definition 1.
A random variable $X$ is said to be $\alpha$-stable if its characteristic function can be written as follows:
@@ 2.1 Definitions and properties
Furthermore, a stable variable is called standard if its scale parameter $\|X\|_{\alpha}$ is equal to 1. In what follows, we will be dealing with $S\alpha S$ random variables.
@@
Depending on the chosen copula linking them, the sum of two Gaussian variables is not necessarily Gaussian. A Gaussian vector is a particular case of the multivariate extension of a Gaussian variable, in which any linear combination of the Gaussian components of the vector is also Gaussian. We can as well define the stable vector as a multivariate generalization of $\alpha$-stable variables in which any linear combination of its components is also $\alpha$-stable [60, Definition 2.1.1]. It is the kind of linear dependence which is used to build a stochastic process with stable increments and we will therefore focus on this case. Two stable variables that constitute a stable vector will be said to be jointly stable.
When $\alpha=2$, the variance of the variable is equal to $2\gamma^{2}$. When $\alpha<2$, $\mathbb{E}(|X|^{p})<\infty$ for any $p\in(0,\alpha)$ and $\mathbb{E}(|X|^{p})=\infty$ for any $p\geq\alpha$ [60, Property 1.2.16]. Therefore, the variance is infinite as soon as $\alpha<2$, but the scale of the variables is appropriately described by the parameter $\gamma$. However, when the variance is infinite, the use of a covariance to describe the dependence between two random variables also becomes impossible. Several dependence measures are proposed in the literature to bypass this limitation [60, Chapter 2]: covariation, Lévy correlation cascade, codifference. Covariation has some practical limitations: it is restricted to $\alpha>1$ and it is based on the spectral measure which can be hardly retrieved with empirical data [49]. Lévy correlation cascade exploits the Poisson process related to the LFSM [23]. Our preference goes to codifference, which is valid whatever $\alpha\in(0,2]$ and which, among the three measures of dependence cited above, is the most related to Definition 1. It is indeed based on the characteristic function of random variables and it can be used for any type of probability distribution, and not only stable ones [40, 7, 56, 72].
Definition 2.
Let $X$ and $Y$ be two random variables and $\Phi_{Z}$ be the characteristic function of any variable $Z$. The codifference between $X$ and $Y$ is defined by
$\operatorname{CD}(X,Y)=-\ln(\Phi_{X}(1))-\ln(\Phi_{Y}(-1))+\ln(\Phi_{X-Y}(1)).$
In the case where $X$ and $Y$ are jointly $S\alpha S$ random variables, with $\alpha\in(0,2]$, the codifference more simply writes
@@
after equation (1). It also has some useful properties that we recall in Proposition 1 [60, Properties 2.10.2-2.10.4].
Proposition 1.
Let $X$ and $Y$ be two jointly $S\alpha S$ random variables, with $\alpha\in(0,2]$. We have the following properties:
  1. 1.
    Symmetry: $\operatorname{CD}(X,Y)=\operatorname{CD}(Y,X)$.
  2. 2.
    Gaussian case: If $\alpha=2$, then $\operatorname{CD}(X,Y)=\operatorname{Cov}(X,Y)$.
  3. 3.
    Independence: If $X$ and $Y$ are independent, then $\operatorname{CD}(X,Y)=0$, whatever $\alpha\in(0,2]$. Reciprocally, $\operatorname{CD}(X,Y)=0$ implies that $X$ and $Y$ are independent, when $\alpha\in(0,1)\cup\{2\}$.
@@
Just as one can generate a Brownian motion by aggregating Gaussian random variables, one can also create a stochastic processes from $\alpha$-stable random variables. It is defined as follows [60, Example 3.1.3].
Definition 3.
A process $(L_{\alpha}(t))_{t\geq 0}$ is an $\alpha$-stable Lévy motion if:
  • (i)
    $L_{\alpha}(0)=0$,
  • (ii)
    $\forall s\leq t,\ L_{\alpha}(t)-L_{\alpha}(s)$ is independent of the $\sigma$-algebra generated by $\{L_{\alpha}(u)|u\leq s\}$,
  • (iii)
    $\forall s\leq t,\ L_{\alpha}(t)-L_{\alpha}(s)$ has the same distribution as $L_{\alpha}(t-s)$, which follows an $\alpha$-stable distribution with scale parameter $|t-s|^{\frac{1}{\alpha}}$.
When $\alpha=2$, the $\alpha$-stable Lévy motion reduces to a Brownian motion of volatility parameter $\sqrt{2}$. Unless we have both $\alpha=1$ and $\beta\neq 0$, the $\alpha$-stable Lévy motion is $1/\alpha$-selfsimilar, meaning that, $\forall c>0$, $\left\{L_{\alpha}(ct)|t\geq 0\right\}$ has the same finite-dimensional distribution as $\left\{c^{1/\alpha}L_{\alpha}(t)|t\geq 0\right\}$. In what follows, we will particularly be interested in the $S\alpha S$ Lévy motion, which is an $\alpha$-stable Lévy motion with $S\alpha S$ increments.
Since we are working with stochastic processes, one needs to define a measure of serial dependence that generalizes autocovariance to non-Gaussian processes. It is the purpose of autocodifference [60, 72].
Definition 4.
Let $(X_{t})_{t\geq 0}$ be any stochastic process. Let $s,t\geq 0$. The autocodifference of the process at times $s$ and $t$ is
$\operatorname{CD}(X_{t},X_{s})=-\ln(\Phi_{X_{t}}(1))-\ln(\Phi_{X_{s}}(-1))+\ln(\Phi_{X_{t}-X_{s}}(1)).$
As a straightforward consequence of Definitions 3 and 4 and equation (1), the autocodifference of an $S\alpha S$ Lévy motion writes [72]
$\operatorname{CD}(L_{\alpha}(t),L_{\alpha}(s))$ $=\|L_{\alpha}(t)\|_{\alpha}^{\alpha}+\|L_{\alpha}(s)\|_{\alpha}^{\alpha}-\|L_{\alpha}(t)-L_{\alpha}(s)\|_{\alpha}^{\alpha}$
$=|t|+|s|-|t-s|$
$=2\min(s,t).$
Since the autocodifference is the codifference applied to observations at two instants of a same stochastic process, we get from the second statement of Proposition 1 that the autocodifference is equal to the autocovariance when considering a Brownian motion of volatility parameter $\sqrt{2}$.
As well as the fBm extends the Brownian motion by modifying its serial dependence structure [46], it is possible to extend the $S\alpha S$ Lévy motion to reproduce specific fractal features. The obtained model is the LFSM. We introduce it by the following integral-based definition [60, 64, 69].
Definition 5.
An LFSM of stability parameter $\alpha\in(0,2]$ and Hurst exponent $H\in(0,1)$ is a stochastic process $(X_{t})_{t\in\mathbb{R}}$ such that, $\forall t\in\mathbb{R}$,
[…]
where $L_{\alpha}(s)$ is an $S\alpha S$ Lévy motion.
The LFSM is a $H$-selfsimilar process with stationary $S\alpha S$ increments [60, Proposition 7.4.2]. The scale parameter of this process is also known to be as follows [60, Proposition 7.4.3][72]:
@@
[…]
The Hurst exponent reflects the dependence that exists between the increments. If $H=1/\alpha$, increments of the LFSM are independent and the process is a standard $S\alpha S$ Lévy motion. If $H>1/\alpha$ (respectively $H<1/\alpha$), increments are positively (resp. negatively) dependent. The further the Hurst exponent is from $1/\alpha$, the strongest the serial dependence will be. More precisely, the serial dependence structure of the LFSM can be described by its autocodifference [72], which we obtain as a consequence of equation (3) and of the stationarity of the increments of $X_{t}$:
@@
2.2 Simulation
Historically, the first simulations of the fBm were based on a discretization of the integral definition of the fBm, that is Definition 5 with $\alpha=2$ [46]. This method only leads to an approximation of a true fBm, for two reasons: the integral is truncated by considering finite bounds, the continuous integral is replaced by a discrete sum. Beside this approximative method, many exact methods have been proposed. Most of these exact simulation methods of the fBm are based on the covariance matrix, like the Cholesky method, or the Davies-Harte and Wood-Chan ones [21, 18].
Unfortunately, these exact methods do not work for an LFSM when $\alpha<2$, because the covariance matrix is not defined. For this reason, the simulation of an LFSM is always approximative. The most popular approach is the Riemann-sum approximation of the stochastic integral representation of the LFSM of Definition 5 [60, Section 7.11][37], with the same two errors as in the fBm case cited above. It is also worth mentioning the existence of an approach that refines the Riemann-sum simulation method by the inclusion of a fast Fourier transform [15, 64].
@@ 2.2 Simulation
2.3 Estimation
Several approaches are proposed in the literature to estimate the two parameters $\alpha$ and $H$ of the LFSM. Estimators based on the wavelet transform have appealing asymptotic properties and are used for the standalone estimation of either $H$ [63, 55] or $\alpha$ [6]. Alternative joint estimation methods can be based on power variations [31, 62], which supposes that one selects a power lower than a known lower bound of $\alpha$. Using an empirical characteristic function is also a natural choice for designing an estimator of both $H$ and $\alpha$ [47, 45]. It indeed follows a widespread method used for estimating the $\alpha$ parameter of a stable distribution [41, 39, 61], among other methods [5], like the one based on empirical quantiles [48].
@@ 2.3 Estimation
Fixing alternatively $\tau$ and $\theta$, on can successively estimate $\alpha$ and $H$.
@@
$\widehat{\Phi}_{Y}:\theta\longmapsto\frac{1}{n}\sum_{i=1}^{n}\cos(Y_{i}).$
@@ 3.1 Discrete-time decomposition of stable processes
When working with Gaussian processes, even with the fBm, the mean and covariance matrix are enough to describe the distribution of a vector of discrete-time observations of this process. This allows, for example, to decompose the components of this vector into sums of independent Gaussian variables for simulation purposes [18], or to forecast future values by a conditional expectation, which we obtain by manipulating matrices [53, 28]. Unfortunately, the stable non-Gaussian case is not a straightforward extension. It has for example been proved that non-trivial continuous-time stable processes do not admit a Karhunen-Loève decomposition, unless $\alpha=2$ [54]. The codifference, introduced in Section 2.1, is in fact not enough to describe all the dependence structure of a stable process. Worse, even the dependence structure of a simple jointly $S\alpha S$ finite-dimensional vector is not totally described by its codifference matrix, that is the matrix containing the codifference between all pairs of components. The missing piece, which totally characterizes the dependence structure, is the spectral measure, which is a measure on the unit sphere $\mathbb{S}^{d-1}$ of $\mathbb{R}^{d}$ [60, Section 2.3]. It appears in the characteristic function of an $S\alpha S$ vector $\mathbf{X}=(X_{1},...,X_{d})^{\prime}$,
@@
where $\Gamma_{\mathbf{X}}$ is the spectral measure, $\boldsymbol{\theta}\in\mathbb{R}^{d}$, and .,$\langle$.,$\rangle$ . , . ⟩ is the scalar product in $\mathbb{R}^{d}$. The codifference between two $S\alpha S$ variables only summarizes in a scalar the dependence structure contained in the spectral measure:
@@
Therefore, two distinct stable vectors, that is with two distinct spectral measures, may have the same codifference matrix. The limitation of the codifference to characterize multivariate stable distributions can also be seen in Proposition 1, where we saw that a zero codifference is a necessary but not sufficient condition for independence.
@@
Let $(X_{t})_{t\in\mathbb{R}}$ be an LFSM and $t\in\mathbb{N}$. Whatever $d\in\mathbb{N}\setminus\{0\}$, if the decomposition proposed in equation (6) exists, with $(X_{t},...,X_{t+d-1})^{\prime}$ having the same codifference matrix as $(\mathcal{T}X_{t},...,\mathcal{T}X_{t+d-1})^{\prime}$, and if the coefficients $a_{t,i,j}$ verify the condition that $a_{t,i,j}>0$ for $i\geq j$ and $a_{t,i^{\prime},j}>a_{t,i,j}$ (respectively $a_{t,i^{\prime},j}<a_{t,i,j}$ and $a_{t,i^{\prime},j}=a_{t,i,j}$) for $j\leq i\leq i^{\prime}$ when $H>1/\alpha$ (resp. $H<1/\alpha$ and $H=1/\alpha$), then the decomposition is unique and is the solution of the following system of $d(d+1)/2$ nonlinear equations:
$\left\{\begin{array}[]{ll}(\mathcal{E}_{i,i})&|a_{t,i,i}|^{\alpha}=K_{\alpha,H}^{\alpha}|t+i|^{\alpha H}-\sum_{j=0}^{i-1}|a_{t,i,j}|^{\alpha}\\ (\mathcal{E}_{i^{\prime},i})&f_{t,i^{\prime},i}(a_{t,i^{\prime},i})=K_{\alpha,H}^{\alpha}\left(|t+i^{\prime}|^{\alpha H}-|i^{\prime}-i|^{\alpha H}\right)-\sum_{j=0}^{i-1}\left(|a_{t,i^{\prime},j}|^{\alpha}-|a_{t,i^{\prime},j}-a_{t,i,j}|^{\alpha}\right)\end{array}\right.$
with equation $(\mathcal{E}_{i,i})$ defined for $i\in\llbracket 0,d-1\rrbracket$, equation $(\mathcal{E}_{i^{\prime},i})$ defined for $(i,i^{\prime})\in\llbracket 0,d-2\rrbracket\times\llbracket i+1,d-1\rrbracket$, and $f_{t,i^{\prime},i}(z)=|z|^{\alpha}-|z-a_{t,i,i}|^{\alpha}$.
@@
The case $\alpha=2$ is also simple and does not require any optimization since $f_{t,i^{\prime},i}(a_{t,i^{\prime},i})=2a_{t,i^{\prime},i}a_{t,i,i}-a_{t,i,i}^{2}$, making it possible to isolate $a_{t,i^{\prime},i}$ in $(\mathcal{E}_{i^{\prime},i})$. This is consistent with the decomposition and forecasting framework for the fBm, which is already well known and only based on matrix manipulations [53, 28].
Theorem 1 focuses on the uniqueness of the solution but not on its existence. However, we have conducted numerical tests with many values of $H$ and $\alpha$ and have always found a solution fulfilling the conditions of the theorem, except for very small values of $\alpha$, as displayed in Figure 3. It is also possible to prove theoretically the existence of the coefficients for small values of $i$. For example, assuming $H>1/\alpha$, in order to prove the existence of $a_{t,1,0}$, we look for $z>a_{t,0,0}=K_{\alpha,H}|t|^{H}$ such that $f_{t,1,0}(z)=K_{\alpha,H}^{\alpha}(|t+1|^{\alpha H}-1)$. We note that $f_{t,1,0}(a_{t,0,0})=K_{\alpha,H}^{\alpha}|t|^{\alpha H}$. The mapping $g:x\mapsto x^{\alpha H}$ is convex because $H>1/\alpha$ and $g(0)=0$, so $g$ is superadditive. Therefore $f_{t,1,0}(a_{t,0,0})\leq K_{\alpha,H}^{\alpha}(|t+1|^{\alpha H}-1)$. We also note that $f_{t,1,0}(K_{\alpha,H}|t+1|^{H})=K_{\alpha,H}^{\alpha}(|t+1|^{\alpha H}-[(t+1)^{H}-t^{H}]^{\alpha})$. Since $H\in(0,1)$, the mapping $h:x\mapsto x^{H}$ is concave and $h(0)=0$ so it is subadditive. Therefore $(t+1)^{H}\leq t^{H}+1$ and $f_{t,1,0}(K_{\alpha,H}|t+1|^{H})\geq K_{\alpha,H}^{\alpha}(|t+1|^{\alpha H}-1)$. Finally, by continuity of $f_{t,1,0}$, we know that a solution $a_{t,1,0}$ to equation $(\mathcal{E}_{1,0})$ exists and that $a_{t,1,0}\in\left[a_{t,0,0},K_{\alpha,H}|t+1|^{H}\right]$.
Figure 3: Top left: Frontier of the pairs $(\alpha,H)$ above which we numerically get the existence of the coefficients $a_{t,i,j}$ respecting the constraints of Theorem 1, for $t=1$ and $i\in\llbracket 0,6\rrbracket$. Three other graphs: coefficients $a_{1,i,j}$ for $i$ between 0 and 6 and $j\in\llbracket 0,i\rrbracket$ (one curve for each $i$), with $(\alpha,H)$ successively equal to $(0.7,0.8)$ (top right), $(1.5,0.8)$ (bottom left), and $(1.5,0.3)$ (bottom right).

### v2

@@ 1 Introduction
One can consider several alternatives to the covariance to build forecasting methods. First, though moments of order $p$ are infinite when $p\leq\alpha$, conditional moments can be finite for higher orders [61, Chapter 5], but it is under a set of assumptions which does not hold when future increments are obtained by adding independent variables to past observations, like in the LFSM [61, Theorem 5.1.3]. The conditional expectation is however finite when another restrictive condition is met, namely when $\alpha>1$ [61, Theorem 5.2.2]. The obtained analytic formula can however not be generalized for a conditioning on a number of lagged increments larger than 1 [60, 35]. A second alternative would consist in describing the dependence structure with copulas. It is however not convenient because it requires an expression for the cumulative distribution function, which, in the case of stable variables, can only be obtained numerically. A third alternative would be to depict the dependence by the mean of the multivariate characteristic function instead of the multivariate cumulative distribution function. Indeed, the spectral measure, which is derived from this multivariate characteristic function, entirely describes the dependence between stable variables. Since it is a function, the spectral measure is an object of infinite dimension, which is thus difficult to estimate from limited observations. Instead, one often considers the codifference [40, 7, 57, 73], which simplifies the dependence structure contained in the spectral measure as the correlation or Kendall rank correlation do with the copula.
In this article, similarly to the Cholesky decomposition of a Gaussian vector, we propose a decomposition of discrete-time observations of an LFSM in independent stable variables. The decomposition has not exactly the same dependence structure as an LFSM but it has the same codifference. From this decomposition, we can propose a forecasting method. The evaluation of this method, first theoretically then on simulations, shows a new way of interpreting the parameters of an LFSM, namely the stability parameter $\alpha$ and the Hurst exponent $H$. Indeed, in addition to the traditional persistent, independent, and antipersistent cases, a new property of the time series appears for very small values of $\alpha$. The method is also compared to a previous study on the fBm using real time series of volatilities or of FX rates. The performance of our forecasting method in this real framework is promising and legitimises the model. Beyond the forecasting performance of the LFSM, we also highlight, through several methods, the non-Gaussian character of volatility series. We thus underline the relevance of a non-Gaussian version of rough volatility models, in which the LFSM can play a pivotal role.
@@
The rest of the article is organized as follows. Section 2 presents the model along with some of its properties and the concept of codifference. In Section 3, we introduce the decomposition of a vector of observations of the LFSM and deduce the forecasting method. A simulation study, with a comparison to the fBm and an autoregressive model, as well as an application to real data, including time series of volatilities, are proposed respectively in Sections 4 and 5. Section 6 concludes.
2 Model description: preliminaries on the LFSM
@@
In this section, we first introduce codifference and LFSM, then we briefly present simulation and estimation methods.
2.1 Codifference of jointly stable variables
Several equivalent definitions of a stable distribution exist [61]. Unfortunately, no analytic expression is available for the probability density function (pdf) but we know the characteristic function of an $\alpha$-stable random variable $X$, which is enough to define this distribution and from which we can get the pdf numerically thanks to an inverse Fourier transform [58, 5]:
@@ 2.1 Codifference of jointly stable variables
In what follows, we will be dealing with $S\alpha S$ random variables, which will be said standard when $\|X\|_{\alpha}=1$.
@@
We are particularly interested in multivariate extensions of $S\alpha S$ variables and in the linear dependence between jointly stable variables, that is between the components of a stable vector [61, Definition 2.1.1]. In general, this dependence cannot be described by the covariance, because the variance itself is infinite, unless $\alpha=2$. Indeed, when $\alpha<2$, $\mathbb{E}(|X|^{p})<\infty$ for any $p\in(0,\alpha)$ and $\mathbb{E}(|X|^{p})=\infty$ for any $p\geq\alpha$ [61, Property 1.2.16].
Several dependence measures are proposed in the literature to bypass this limitation [61, Chapter 2]: covariation, Lévy correlation cascade, codifference. Covariation has some practical limitations: it is restricted to $\alpha>1$ and it is based on the spectral measure which can be hardly retrieved with empirical data [49]. Lévy correlation cascade exploits the Poisson process related to the LFSM [23]. Our preference goes to codifference, which is valid whatever $\alpha\in(0,2]$ and which, among the three measures of dependence cited above, is the most related to the only available characterization of stable variables. It is indeed based on the characteristic function of random variables and it can be used for any type of probability distribution, and not only stable ones [40, 7, 57, 73].
The codifference between two random variables $X$ and $Y$ is defined by
$\operatorname{CD}(X,Y)=-\ln(\Phi_{X}(1))-\ln(\Phi_{Y}(-1))+\ln(\Phi_{X-Y}(1)),$
where $\Phi_{Z}$ is the characteristic function of any variable $Z$. In the case where $X$ and $Y$ are jointly $S\alpha S$ random variables, with $\alpha\in(0,2]$, the codifference more simply writes
@@
after equation (1). For jointly $S\alpha S$ random variables $X$ and $Y$, codifference also has some useful properties: it is symmetric in $X$ and $Y$; when $\alpha=2$ it is equal to covariance; it is equal to zero when $X$ and $Y$ are independent of each other, regardless of $\alpha\in(0,2]$, the converse being true when $\alpha\in(0,1)\cup\{2\}$ [61, Properties 2.10.2-2.10.4].
@@
2.2 The LFSM
The $S\alpha S$ Lévy motion is the extension of a Brownian motion with an $S\alpha S$ distribution instead of the Gaussian one [61, Example 3.1.3]. As well as the fBm extends the Brownian motion by modifying its serial dependence structure [46], it is possible to extend the $S\alpha S$ Lévy motion to reproduce specific fractal features. The obtained process is the LFSM $(X_{t})_{t\in\mathbb{R}}$, of stability parameter $\alpha\in(0,2]$ and Hurst exponent $H\in(0,1)$, defined for all $t\in\mathbb{R}$ by
$X_{t}=\int_{\mathbb{R}}\left[(t-s)_{+}^{H-\frac{1}{\alpha}}-(-s)_{+}^{H-\frac{1}{\alpha}}\right]dL_{\alpha}(s),$ (3)
where $L_{\alpha}(s)$ is an $S\alpha S$ Lévy motion [61, 65, 70].
The LFSM is thus a $H$-selfsimilar process with stationary $S\alpha S$ increments [61, Proposition 7.4.2]. The scale parameter of this process is also known to be as follows [61, Proposition 7.4.3][73]:
@@ 2.2 The LFSM
$K_{\alpha,H}=\left(\int_{\mathbb{R}}{\left|(1-s)_{+}^{H-\frac{1}{\alpha}}-(-s)_{+}^{H-\frac{1}{\alpha}}\right|^{\alpha}ds}\right)^{\frac{1}{\alpha}}.$
The Hurst exponent reflects the dependence that exists between increments. If $H=1/\alpha$, the increments of the LFSM are independent and the process is a standard $S\alpha S$ Lévy motion. If $H>1/\alpha$ (respectively $H<1/\alpha$), the increments are positively (resp. negatively) dependent. The further the Hurst exponent is from $1/\alpha$, the strongest the serial dependence will be. More precisely, the serial dependence structure of the LFSM can be described by its autocodifference [73], which we obtain as a consequence of equation (4) and of the stationarity of the increments of $X_{t}$:
@@
When $H=1/\alpha$ the above formula reduces to $\operatorname{CD}(X_{t},X_{s})=2\min(s,t)$.
2.3 Simulation of the LFSM
Historically, the first simulations of the fBm were based on a discretization of the integral definition of the fBm, that is equation (3) with $\alpha=2$ [46]. This method only leads to an approximation of a true fBm, for two reasons: the integral is truncated by considering finite bounds, the continuous integral is replaced by a discrete sum. Beside this approximative method, many exact methods have been proposed. Most of these exact simulation methods of the fBm are based on the covariance matrix, like the Cholesky method, or the Davies-Harte and Wood-Chan ones [21, 18].
Unfortunately, these exact methods do not work for an LFSM when $\alpha<2$, because the covariance matrix is not defined. For this reason, the simulation of an LFSM is always approximative. The most popular approach is the Riemann-sum approximation of the stochastic integral representation of the LFSM of equation (3) [61, Section 7.11][37], with the same two errors as in the fBm case cited above. It is also worth mentioning the existence of an approach that refines the Riemann-sum simulation method by the inclusion of a fast Fourier transform [15, 65].
@@ 2.3 Simulation of the LFSM
2.4 Estimation of the LFSM
Several approaches are proposed in the literature to estimate the two parameters $\alpha$ and $H$ of the LFSM. Estimators based on the wavelet transform have appealing asymptotic properties and are used for the standalone estimation of either $H$ [64, 56] or $\alpha$ [6]. Alternative joint estimation methods can be based on power variations [31, 63], which supposes that one selects a power lower than a known lower bound of $\alpha$. Using an empirical characteristic function is also a natural choice for designing an estimator of both $H$ and $\alpha$ [47, 45]. It indeed follows a widespread method used for estimating the $\alpha$ parameter of a stable distribution [41, 39, 62], among other methods [53, 5], like the one based on empirical quantiles [48].
@@ 2.4 Estimation of the LFSM
Fixing alternatively $\tau$ and $\theta$, one can successively estimate $\alpha$ and $H$.
@@
$\widehat{\Phi}_{Y}:\theta\longmapsto\frac{1}{n}\sum_{i=1}^{n}\cos(\theta Y_{i}).$
@@ 3.1 Discrete-time decomposition of stable processes
When working with Gaussian processes, even with the fBm, the mean and covariance matrix are enough to describe the distribution of a vector of discrete-time observations of this process. This allows, for example, to decompose the components of this vector into sums of independent Gaussian variables for simulation purposes [18], or to forecast future values by a conditional expectation, which we obtain by manipulating matrices [54, 28]. Unfortunately, the stable non-Gaussian case is not a straightforward extension. It has for example been proved that non-trivial continuous-time stable processes do not admit a Karhunen-Loève decomposition, unless $\alpha=2$ [55]. The codifference, introduced in Section 2.1, is in fact not enough to describe all the dependence structure of a stable process. Worse, even the dependence structure of a simple jointly $S\alpha S$ finite-dimensional vector is not totally described by its codifference matrix, that is the matrix containing the codifference between all pairs of components. The missing piece, which totally characterizes the dependence structure, is the spectral measure, which is a measure on the unit sphere $\mathbb{S}^{d-1}$ of $\mathbb{R}^{d}$ [61, Section 2.3]. It appears in the characteristic function of an $S\alpha S$ vector $\mathbf{X}=(X_{1},...,X_{d})^{\prime}$,
@@
where $\Gamma_{\mathbf{X}}$ is the spectral measure, $\boldsymbol{\theta}\in\mathbb{R}^{d}$, and $\langle.,.\rangle$ is the scalar product in $\mathbb{R}^{d}$. The codifference between two $S\alpha S$ variables only summarizes in a scalar the dependence structure contained in the spectral measure:
@@
Therefore, two distinct stable vectors, that is with two distinct spectral measures, may have the same codifference matrix. The limitation of the codifference to characterize multivariate stable distributions can also be seen in Section 2.1, where we saw that a zero codifference is a necessary but not sufficient condition for independence.
@@
Let $(X_{t})_{t\in\mathbb{R}}$ be an LFSM and $t\in\mathbb{N}$. Whatever $d\in\mathbb{N}\setminus\{0\}$, if the decomposition proposed in equation (7) exists, with $(X_{t},...,X_{t+d-1})^{\prime}$ having the same codifference matrix as $(\mathcal{T}X_{t},...,\mathcal{T}X_{t+d-1})^{\prime}$, and if the coefficients $a_{t,i,j}$ verify the condition that $a_{t,i,j}>0$ for $i\geq j$ and $a_{t,\ell,j}>a_{t,i,j}$ (respectively $a_{t,\ell,j}<a_{t,i,j}$ and $a_{t,\ell,j}=a_{t,i,j}$) for $j\leq i\leq\ell$ when $H>1/\alpha$ (resp. $H<1/\alpha$ and $H=1/\alpha$), then the decomposition is unique and is the solution of the following system of $d(d+1)/2$ nonlinear equations:
$\left\{\begin{array}[]{ll}(\mathcal{E}_{i,i})&|a_{t,i,i}|^{\alpha}=K_{\alpha,H}^{\alpha}|t+i|^{\alpha H}-\sum_{j=0}^{i-1}|a_{t,i,j}|^{\alpha}\\ (\mathcal{E}_{\ell,i})&f_{t,i}(a_{t,\ell,i})=K_{\alpha,H}^{\alpha}\left(|t+\ell|^{\alpha H}-|\ell-i|^{\alpha H}\right)-\sum_{j=0}^{i-1}\left(|a_{t,\ell,j}|^{\alpha}-|a_{t,\ell,j}-a_{t,i,j}|^{\alpha}\right)\end{array}\right$.
with equation $(\mathcal{E}_{i,i})$ defined for $i\in\llbracket 0,d-1\rrbracket$, equation $(\mathcal{E}_{\ell,i})$ defined for $(i,\ell)\in\llbracket 0,d-2\rrbracket\times\llbracket i+1,d-1\rrbracket$, and $f_{t,i}(z)=|z|^{\alpha}-|z-a_{t,i,i}|^{\alpha}$.
@@
The case $\alpha=2$ is also simple and does not require any optimization since $f_{t,i}(a_{t,\ell,i})=2a_{t,\ell,i}a_{t,i,i}-a_{t,i,i}^{2}$, making it possible to isolate $a_{t,\ell,i}$ in $(\mathcal{E}_{\ell,i})$. This is consistent with the decomposition and forecasting framework for the fBm, which is already well known and only based on matrix manipulations [54, 28].
Theorem 1 focuses on the uniqueness of the solution but not on its existence. However, we have conducted numerical tests with many values of $H$ and $\alpha$ and have always found a solution fulfilling the conditions of the theorem, except for very small values of $\alpha$, much smaller than the ones estimated in the empirical study of Section 5, as displayed in Figure 3. It is also possible to prove theoretically the existence of the coefficients for small values of $i$. For example, assuming $H>1/\alpha$, in order to prove the existence of $a_{t,1,0}$, we look for $z>a_{t,0,0}=K_{\alpha,H}|t|^{H}$ such that $f_{t,0}(z)=K_{\alpha,H}^{\alpha}(|t+1|^{\alpha H}-1)$. We note that $f_{t,0}(a_{t,0,0})=K_{\alpha,H}^{\alpha}|t|^{\alpha H}$. The mapping $g:x\mapsto x^{\alpha H}$ is convex because $H>1/\alpha$ and $g(0)=0$, so $g$ is superadditive. Therefore $f_{t,0}(a_{t,0,0})\leq K_{\alpha,H}^{\alpha}(|t+1|^{\alpha H}-1)$. We also note that $f_{t,0}(K_{\alpha,H}|t+1|^{H})=K_{\alpha,H}^{\alpha}(|t+1|^{\alpha H}-[(t+1)^{H}-t^{H}]^{\alpha})$. Since $H\in(0,1)$, the mapping $h:x\mapsto x^{H}$ is concave and $h(0)=0$ so it is subadditive. Therefore $(t+1)^{H}\leq t^{H}+1$ and $f_{t,0}(K_{\alpha,H}|t+1|^{H})\geq K_{\alpha,H}^{\alpha}(|t+1|^{\alpha H}-1)$. Finally, by continuity of $f_{t,0}$, we know that a solution $a_{t,1,0}$ to equation $(\mathcal{E}_{1,0})$ exists and that $a_{t,1,0}\in\left[a_{t,0,0},K_{\alpha,H}|t+1|^{H}\right]$.
Figure 3: Top left: Frontier of the pairs $(\alpha,H)$ above which we numerically get the existence of the coefficients $a_{t,i,j}$ respecting the constraints of Theorem 1, for $t=1$ and $i\in\llbracket 0,6\rrbracket$. Three other graphs: coefficients $a_{1,i,j}$ for $i$ between 0 and 6 and $j\in\llbracket 0,i\rrbracket$ (one curve for each ii), with $(\alpha,H)$ successively equal to $(0.7,0.8)$ (top right), $(1.5,0.8)$ (bottom left), and $(1.5,0.3)$ (bottom right).